the ratio of length to width of a rectangle is 5 to 2 . what are the dimensions of the rectangle if its perimeter is 98cm . please help me solve step by step , thank you
Answer by rapaljer(4671)

Let 2x= width
5x= length

Remember that 2W + 2L = P
2(2x)+2(5x)=98
4x+10x=98
14x=98
x=98/14=7 cm
2x = 2*2 = 14 cm = WIDTH
5x = 5*7 = 35 cm = LENGTH

Check:
2*14+2*35
28+70= 98 cm It checks!!

For additional help with all kinds of WORD PROBLEMS, especially Perimeter problems, please see my own website. Just click on my tutor name "rapaljer" anywhere in algebra.com. A link there will connect you to my website. On my Homepage, look for the link "Basic, Intermediate, and College Algebra: One Step at a Time." Choose "Basic Algebra", and look in "Chapter 1" for the topics on Word Problems, Sections 1.09 and 1.10". Also see solutions to these kinds of problems on my "MATH IN LIVING COLOR" pages that corrspond to these pages. Most people tell me that my explanations are a lot easier to understand than their expensive textbooks--and MINE ARE ALL FREE!!! Imgaine that!! If you like my explanations, all I ask is that you recommend them to friends and family!!

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus